# Phase diagram of the quantum spin-1/2 Heisenberg-$\Gamma$ model on a frustrated zigzag chain

### v1

@@ Abstract
We investigate the quantum spin-1/2 zigzag chain with frustrated $J_{1}$-$J_{2}$ Heisenberg interactions, incorporating additional off-diagonal exchange interactions known as the $\Gamma$-term, both with and without an applied magnetic field. Based on the density-matrix renormalization group calculation, we map out the ground state phase diagram that shows a variety of magnetic and nonmagnetic phases including multicritical points and several exactly solvable points. Upon introducing a finite $\Gamma$-term, we observe the persistent dimer singlet state of the $J_{1}$-$J_{2}$ Heisenberg model, sustaining a nonzero spin gap, while also giving rise to a new gapless branch of nonmagnetic excitation. This gapless mode induces robust nematic fluctuations manifesting in its long-ranged correlations, however, the true condensation does not occur until the $\Gamma$-term becomes comparable with the Heisenberg term. There, the nonmagnetic phases transform into Ising-type ferromagnetic or antiferromagnetic long-range orders that arise from the $\Gamma$-term spontaneously selecting magnetic easy axes. Its orientations dictate the type of magnetic order under geometric frustration effects as predicted by Landau’s mean-field theory. These theoretical findings provide insights into the exotic low-temperature phase observed in YbCuS2, characterized by gapless excitations and seemingly nonmagnetic behavior accompanied by incommensurate correlations.
@@ I Introduction
Exploring interesting quantum-disordered phases in materials stands by now as a major challenge in condensed matter physics. Longstanding intense investigations into triangular and kagome quantum spin liquidsAnderson (1973, 1987); Zhou et al. (2017) were recently spurred by the discovery of Kitaev spin liquids featuring unique quasiparticle excitationsKitaev (2006), and it motivated the detailed examination of new types of quantum anisotropic exchange interactions in 4d, 5d, and 4f insulating magnetsJackeli and Khaliullin (2009); Rau and Gingras (2018); Rau et al. (2016). These interactions are influenced by strong spin-orbit coupling, moderate crystal field effects, and electron correlations, which play a crucial role in the emergence of Kitaev and $\Gamma$ terms alongside the previously studied Dzyaloshinskii-Moriya and ring exchange interactions.
@@
Spin liquids are, however, not the sole focus of nonmagnetic disordered phases in quantum magnets. There are valence-bond solids in the spin-1 chain known as symmetry-protected topological phasePollmann et al. (2010), valence-bond crystals based on singlets in Shastley-Sutherland modelMiyahara and Ueda (1999); Nomura et al. (2023), and spin nematic phases or a quadrupolar order triggered by the condensation of two-magnon bound state for $S=1$ modelsPenc and Läuchli (2011); Chubukov (1991); Harada and Kawashima (2002) and $S=1/2$ spin laddersHikihara et al. (2008); Läuchli et al. (2006). Notably, some of these phases break lattice symmetry while suppressing magnetic orderings, making them experimentally more accessible compared to spin liquids. Fortunately, there exist material platforms that host these phases, such as $\kappa$-ET2Cu2(CN)3Shimizu et al. (2003) and ZnCu3(OH)6Cl2Mendels and Bert (2015); Norman (2016) for spin liquids, NENP for Haldane chainLu et al. (1991), and SrCu2(BO3)2Kageyama et al. (1999) for orthogonal dimer phases, providing crucial insights into the nature of these intriguing phases governed by strong quantum fluctuations and correlations.
@@
Despite these advancements, the role of off-diagonal symmetric $\Gamma$ terms, which have recently been observed in Yb and other 4f-based magnets, remains largely unexplored Chaloupka et al. (2010); Rau et al. (2016). Initially discussed as secondary terms in Kitaev magnets, the $\Gamma$ term, in conjunction with the Heisenberg interaction, destabilizes Kitaev spin liquids in two dimensionsRau et al. (2014). In the one-dimensional analog known as the Kitaev-Heisenberg-$\Gamma$ chain Yang et al. (2020), the $\Gamma$ term significantly alters the ground state phase diagram, leading to SU(2) symmetric points and magnetic orderings with spontaneously oriented easy axes.
@@
This paper elucidates the role of the $\Gamma$ term in geometrically frustrated zigzag Heisenberg spin-1/2 chain, whose potential platform is the 4f insulating magnet, YbCuS2. Previously, we have microscopically derived the quantum spin model for this material based on the $\Gamma_{6}$ Kramers doublet of Yb ions that form a zigzag chain, revealing nearly isotropic $J_{1}\sim J_{2}$ Heisenberg interactions and small but finite $\Gamma$-type exchange couplings Saito et al. (2024).
@@
Figure 1: (a) Ground-state phase diagram of a spin-1/2 zigzag spin chain shown in the right panel, where the $z$-axis parallel to the legs and the $y$-axis perpendicular to the triangular plane. We take $J_{1}=\cos\phi\cos\theta,J_{2}=\sin\phi\cos\theta,\Gamma_{1}=\cos\phi\sin\theta$ and $\Gamma_{2}=\sin\phi\sin\theta$. Three nonmagnetic phases are denoted as the nematic-singlet dimer (ND) phase, the Tomonaga-Luttinger liquid (TLL) phase, and vector chiral (VC) phase. Magnetically long-range ordered phases are ferromagnetic (F), antiferromagnetic (AF) UD, or UUDD phases. Four circles represent the exact solution points (tricritical Lifshitz point, nematic product state, RVB state, fully polarized ferromagnet), and the bold lines (MG-line and $\theta=0$ line) are the exact solution lines. Pink and dark phase boundaries are determined by the change of magnetization and the energy crossing, respectively, and the VC and TLL-ND phases are from the scaling and vector chiral correlations, respectively. (b) Characteristics of phases in terms of magnetic order, correlation functions, and spin gap, to be detailed in Figs. 2-5.
Figure 2: (a) Energy obtained by DMRG calculation and (b) $\langle S^{z}_{i}S^{z}_{i+1}-S^{x}_{i}S^{x}_{i+1}\rangle$ obtained by the exact diagonalization with OBC as functions of $\phi$. The kinks in (a) and crossings in (b) are used to determine the magnetic and nonmagnetic phase boundaries, respectively. Broken lines in panel (a) are shown as a guide to determine one of the phase boundaries.
Figure 3: Uniform and staggered magnetizations in a period of two or four sites, $S^{\alpha}_{i}+S^{\alpha}_{i+1}+S^{\alpha}_{i+2}+S^{\alpha}_{i+3}$, $S^{\alpha}_{i}-S^{\alpha}_{i+1}+S^{\alpha}_{i+2}-S^{\alpha}_{i+3}$, and $S^{\alpha}_{i}\pm S^{\alpha}_{i+1}-S^{\alpha}_{i+2}\mp S^{\alpha}_{i+3}$, for (a) $\theta/\pi=0.1$ in $\alpha=y^{\prime}$-direction and (b) $\theta/\pi=0.4$ in $x^{\prime}$ and $y$’-directions as functions of $\phi$. (c) Vector chiral correlation function $\overline{|\langle\kappa^{(n)}_{i}\kappa^{(n^{\prime})}_{j}\rangle|}$ averaged over $|i-j|=58-61$ as function $\phi$. Gray region has relatively weak signal of chiral correlation. (d) Peak position $q$ of the structure factor ${\cal S}^{\alpha\alpha^{\prime}}(q)$ for $\alpha\alpha^{\prime}=xx,zz,xy$ as a function of $\phi$. We set $\theta/\pi=0.125$ for (c,d).
@@ II.1 Model Hamiltonian
where $J_{\eta}$ and $\Gamma_{\eta}$ are the Heisenberg and anisotropic exchange interactions between nearest ($\eta=1$) and next nearest ($\eta=2$) spins. We consider both the antiferromagnetic(AF) and ferromagnetic(F) couplings of $J_{\eta}$ and set
@@
where dividing the parameter range into four, $\phi=[0:\pi/2],[\pi/2:\pi],[\pi:3\pi/2],[3\pi/2:2\pi]$ correspond to AF-AF, F-AF, F-F, AF-F interactions of $J_{1}$-$J_{2}$, respectively. The sign of $\Gamma_{\eta}$ can be converted by the local unitary transformation and does not influence the physical state. The spin quantization axis $z$ is taken parallel to the chain (see Fig. 1).
@@
The model at $\Gamma_{\eta}=0$ corresponds to the zigzag Heisenberg spin chain, the phase diagram of which has been extensively studied previously. In the antiferromagnetic-antiferromagnetic (AF-AF) case where $J_{1}$ and $J_{2}$ are both positive, there exists a transition from a Tomonaga-Luttinger liquid (TLL) phase to a dimer singlet phase characterized by a finite spin gap Majumdar and Ghosh (1969a, b); Haldane (1982). This transition occurs at a critical point $(J_{2}/J_{1})_{c}\approx 0.2411$ ($\phi\sim 0.148$) Okamoto and Nomura (1992); Eggert (1996). Within the dimer singlet phase, a Lifshitz point $(J_{2}/J_{1})_{L}\approx 0.5206$ delineates two distinct regions: one with commensurate $q=\pi$ short-range magnetic correlations and another with incommensurate $q<\pi$ correlations at lower and higher $J_{2}/J_{1}$ values, respectively Hikihara et al. (2001); Bursill et al. (1995). There is also an ongoing debate regarding a potential transition from the dimer phase to a gapless phase around $J_{2}/J_{1}\approx 2.2$, identified through level crossing experiments Kumar et al. (2015), exact diagonalization and density matrix renormalization group (DMRG) studies Soos et al. (2016). This observation contradicts field theory predictions, which suggest a finite gap $\Delta\sim\exp(-(J_{2}/J_{1})^{\eta})$ with $\eta=1$ White and Affleck (1996) or $\eta=2/3$ Itoi and Qin (2001) for large $J_{2}/J_{1}$ values.
@@
In the F-AF case, a Haldane dimer phase exists where ferromagnetic spin pairs form $S=1$ states for $J_{2}/J_{1}<-1/4$ Furukawa et al. (2012); Agrapidis et al. (2019). At the critical point $J_{2}/J_{1}=-1/4$, an exactly solvable resonating valence bond (RVB) state appears Hamada et al. (1988), leading to a transition to a phase with ferromagnetic (F) long-range order. Similar to the AF-AF case, the existence of a gap at large $J_{2}/|J_{1}|$ remains a subject of controversy Kumar et al. (2015); Agrapidis et al. (2019).
For the F-F and AF-F regions, there is no frustration and the ferromagnetic and TLL phases are realized, respectively.
@@
The zigzag XXZ chain exhibits richer variety of phases because of the broken SU(2) symmetry, including dimer singlet, spin fluid, ferromagnetic, antiferromagnetic phasesSomma and Aligia (2001); Plekhanov et al. (2010); Hirata and Nomura (2000); Gerhardt et al. (1998); Jafari and Langari (2007), and long-range order of vector chirality at large Ising anisotropyHikihara et al. (2001). Similarly, a zigzag Heisenberg ladder in an applied magnetic field also displays comparably a Tomonaga-Luttinger liquid (TLL), a $1/3$-magnetization plateau, spin-density-wave, vector chiral, and fully polarized phases Okunishi and Tonegawa (2003); Okunishi (2008); Hikihara et al. (2010, 2008); Sudan et al. (2009).

### v2

@@ Abstract
We investigate the quantum spin-1/2 zigzag chain with frustrated $J_{1}$-$J_{2}$ Heisenberg interactions, incorporating additional off-diagonal exchange interactions known as the $\Gamma$ term, both with and without an applied magnetic field. Based on the density-matrix renormalization group calculation, we map out the ground state phase diagram that shows a variety of magnetic and nonmagnetic phases including multicritical points and several exactly solvable points. Upon introducing a finite $\Gamma$ term, we observe the persistent dimer singlet state of the $J_{1}$-$J_{2}$ Heisenberg model, sustaining a nonzero spin gap, while also giving rise to a gapless nonmagnetic excitation, manifesting in the substantial zero-energy peak in the nematic dynamical structure factor. This gapless peak-mode remaining almost as a fluctuation to the ground state, induces dilute but robust concentration of nematicity on top of singlets on dimers, which we call the nematic singlet-dimer phase. When the whole nematic excited mode condenses and replaces the singlet, the nematic-dimer phase transforms to the Ising-type ferromagnetic or antiferromagnetic long-range orders that arise from the $\Gamma$ term spontaneously selecting magnetic easy axes. Its orientations dictate the type of magnetic order under geometric frustration effects as predicted by Landau’s mean-field theory. These theoretical findings provide insights into the exotic low-temperature phase observed in YbCuS2, characterized by gapless excitations and seemingly nonmagnetic behavior accompanied by incommensurate correlations.
@@ I Introduction
Exploring interesting quantum-disordered phases in materials stands by now as a major challenge in condensed matter physics. Longstanding intense investigations into triangular and kagome quantum spin liquidsAnderson (1973, 1987); Zhou et al. (2017) were recently spurred by the discovery of Kitaev spin liquids featuring Majorana quasiparticle excitationsKitaev (2006), and it motivated the detailed examination of new types of quantum anisotropic exchange interactions in 4d, 5d, and 4f insulating magnetsJackeli and Khaliullin (2009); Chaloupka et al. (2010); Rau and Gingras (2018); Rau et al. (2016). These interactions are influenced by strong spin-orbit coupling, moderate crystal field effects, and electron correlations, which play a crucial role in the emergence of Kitaev and $\Gamma$ terms alongside the previously studied Dzyaloshinskii-Moriya and ring exchange interactions.
@@
Spin liquids are, however, not the sole focus of nonmagnetic disordered phases in quantum magnets. There are valence-bond solids in the spin-1 chain known as symmetry-protected topological phasePollmann et al. (2010), valence-bond crystals based on singlets in Shastley-Sutherland modelMiyahara and Ueda (1999); Nomura et al. (2023), and spin nematic phases or a quadrupolar order triggered by the condensation of two-magnon bound state for $S=1$ modelsPenc and Läuchli (2011); Chubukov (1991); Harada and Kawashima (2002) and $S=1/2$ or $S=1$ spin laddersHikihara et al. (2008); Läuchli et al. (2006). Notably, some of these phases break lattice symmetry while suppressing magnetic orderings, making them experimentally more accessible compared to spin liquids. Fortunately, there exist material platforms that host these phases, such as $\kappa$-ET2Cu2(CN)3Shimizu et al. (2003) and ZnCu3(OH)6Cl2Mendels and Bert (2015); Norman (2016), for spin liquids, NENP for Haldane chainLu et al. (1991), and SrCu2(BO3)2Kageyama et al. (1999) for orthogonal dimer phases, providing crucial insights into the nature of these intriguing phases governed by strong quantum fluctuations and correlations.
@@
Despite these advancements, the role of off-diagonal symmetric $\Gamma$ terms, which have recently been observed in Yb and other 4f-based magnets, remains largely unexplored Rau and Gingras (2018); Rau et al. (2016). Initially discussed as secondary terms in Kitaev magnets, the $\Gamma$ term, in conjunction with the Heisenberg interaction, destabilizes Kitaev spin liquids in two dimensionsRau et al. (2014). In the one-dimensional analog known as the Kitaev-Heisenberg-$\Gamma$ chain Yang et al. (2020), the $\Gamma$ term significantly alters the ground state phase diagram, leading to SU(2) symmetric points and magnetic orderings with spontaneously oriented easy axes.
@@
This paper elucidates the role of the $\Gamma$ term in geometrically frustrated zigzag Heisenberg spin-1/2 chain, whose potential platform is the 4f insulating magnet, YbCuS2. Previously, we have microscopically derived the quantum spin model for this material based on the $\Gamma_{6}$ Kramers doublet of Yb ions that forms a zigzag chain, revealing nearly isotropic $J_{1}\sim J_{2}$ Heisenberg interactions and small but finite $\Gamma$-type exchange couplings Saito et al. (2024).
@@
Figure 1: (a) Ground-state phase diagram of a spin-1/2 zigzag spin chain shown in the right panel, where the $z$-axis parallel to the legs and the $y$-axis perpendicular to the triangular plane. We take $J_{1}=\cos\phi\cos\theta,J_{2}=\sin\phi\cos\theta,\Gamma_{1}=\cos\phi\sin\theta$ and $\Gamma_{2}=\sin\phi\sin\theta$. Three nonmagnetic phases are denoted as the nematic-singlet dimer (ND) phase, the Tomonaga-Luttinger liquid (TLL) phase, and vector chiral (VC) phase. Magnetically long-range ordered phases are ferromagnetic (FM), antiferromagnetic (AFM) UD, or UUDD phases. Four circles represent the exact solution points (tricritical Lifshitz point, nematic product state, RVB state, fully polarized ferromagnet), and the bold lines (Majumdar-Ghosh line and $\theta=0$ line) are the exact solution lines. Pink and dark phase boundaries are determined by the change of magnetization and the energy crossing, respectively, and the VC and TLL-ND phases are from the scaling and vector chiral correlations, respectively. (b) Characteristics of phases in terms of magnetic order, correlation functions, and spin gap, to be detailed in Figs. 2-5.
@@ II.1 Model Hamiltonian
where $J_{\eta}$ and $\Gamma_{\eta}$ are the Heisenberg and anisotropic exchange interactions between nearest ($\eta=1$) and next nearest ($\eta=2$) spins. We consider both the antiferromagnetic(AFM) and ferromagnetic(FM) couplings of $J_{\eta}$ and set
@@
where dividing the parameter range into four, $\phi=[0:\pi/2],[\pi/2:\pi],[\pi:3\pi/2],[3\pi/2:2\pi]$ correspond to AFM-AFM, FM-AFM, FM-FM, AFM-FM interactions of $J_{1}$-$J_{2}$, respectively. The sign of $\Gamma_{\eta}$ can be converted by the local unitary transformation and does not influence the physical state. The spin quantization axis $z$ is taken parallel to the chain (see Fig. 1).
@@
The model at $\Gamma_{\eta}=0$ corresponds to the zigzag Heisenberg spin chain, the phase diagram of which has been extensively studied previously. In the AFM-AFM case where $J_{1}$ and $J_{2}$ are both positive, there exists a transition from a Tomonaga-Luttinger liquid (TLL) phase to a dimer singlet phase characterized by a finite spin gap Majumdar and Ghosh (1969a, b); Haldane (1982). This transition occurs at a critical point $(J_{2}/J_{1})_{c}\approx 0.2411$ ($\phi\approx 0.075\pi$) Okamoto and Nomura (1992); Eggert (1996). Within the dimer singlet phase, a Lifshitz point $(J_{2}/J_{1})_{L}\approx 0.5206$ ($\phi\approx 0.153\pi$) delineates two distinct regions: one with commensurate $q=\pi$ short-range magnetic correlations and another with incommensurate $q<\pi$ correlations at lower and higher $J_{2}/J_{1}$ values, respectively Hikihara et al. (2001); Bursill et al. (1995). There is also an ongoing debate regarding a potential transition from the dimer phase to a gapless phase around $J_{2}/J_{1}\approx 2.2$ ($\phi\approx 0.36\pi$), identified through level crossing experiments Kumar et al. (2015), exact diagonalization and density matrix renormalization group (DMRG) studies Soos et al. (2016). This observation contradicts field theory predictions, which suggest a finite gap $\Delta\sim\exp(-(J_{2}/J_{1})^{\eta})$ with $\eta=1$ White and Affleck (1996) or $\eta=2/3$ Itoi and Qin (2001) for large $J_{2}/J_{1}$ values. We see shortly in our phase diagram in Fig. 1(a) that the phase boundary has a kink which extrapolates to $J_{2}/J_{1}\approx 2.2$, indicating that the phase changes its nature at $J_{2}/J_{1}\gtrsim 2.2$. Whereas, it is numerically difficult to conclude whether there is a finite but exponentially small spin gap or not.
@@
In the FM-AFM case, a Haldane dimer phase exists where ferromagnetic spin pairs form $S=1$ states for $J_{2}/J_{1}<-1/4$ ($0.922\pi<\phi<\pi$) Furukawa et al. (2012); Agrapidis et al. (2019). At the critical point $J_{2}/J_{1}=-1/4$ ($\phi=0.922\pi$), an exactly solvable resonating valence bond (RVB) state appears Hamada et al. (1988), leading to a transition to a phase with ferromagnetic (FM) long-range order. Similar to the AFM-AFM case, the existence of a gap at large $J_{2}/|J_{1}|$ remains a subject of controversy Kumar et al. (2015); Agrapidis et al. (2019).
For the FM-FM and AFM-FM regions, there is no frustration and the ferromagnetic and TLL phases are realized, respectively.
@@
The zigzag XXZ chain exhibits a richer variety of phases because of the broken SU(2) symmetry, including dimer singlet, spin fluid, ferromagnetic, antiferromagnetic phasesSomma and Aligia (2001); Plekhanov et al. (2010); Hirata and Nomura (2000); Gerhardt et al. (1998); Jafari and Langari (2007), and long-range order of vector chirality at large Ising anisotropyHikihara et al. (2001). Similarly, a zigzag Heisenberg ladder in an applied magnetic field also displays comparably a Tomonaga-Luttinger liquid (TLL), a $1/3$-magnetization plateau, spin-density-wave, vector chiral, and fully polarized phases Okunishi and Tonegawa (2003); Okunishi (2008); Hikihara et al. (2010, 2008); Sudan et al. (2009).